# Detection of pairwise kSZ effect with DESI galaxy groups and Planck in Fourier space

### 07 Jan 2024 v1

@@ Abstract
We report a $\sim 5.2\sigma$ detection of the kinetic Sunyaev–Zel’dovich (kSZ) effect in Fourier space, by combining the DESI galaxy groups and the Planck data. We use the density-weighted pairwise kSZ power spectrum as the summary statistic, and the detailed procedure of its measurement is presented in this paper. Meanwhile, we analyze the redshift space group density power spectrum to constrain its bias parameters and photo-z uncertainties. These best fitted parameters are substituted to a non-linear kSZ model, and we fit the measured kSZ power spectrum with this model to constrain the group optical depth $\bar{\tau}$. Selected by a varying lower mass threshold $M_{\rm th}$, the galaxy group catalogs with different median masses ($\tilde{M}$) are constructed from the DR9 data of the DESI Legacy Imaging Surveys. $\tilde{M}$ spans a wide range of $\sim 10^{13}-10^{14}{\rm M}_{\odot}/h$ and the heaviest $\tilde{M}\sim 10^{14}{\rm M}_{\odot}/h$ is larger than those of most other kSZ detections. When the aperture photometric filter radius $\theta_{\rm AP}$ is set to be $4.2$ arcmin, the $\tilde{M}=1.75\times 10^{13}{\rm M}_{\odot}/h$ group sample at the median redshift $\tilde{z}=0.64$ has the highest kSZ detection ${\rm S/N}=5.2$. By fitting $\bar{\tau}$s from various samples against their $\tilde{M}$s, we obtain a linear $\log\bar{\tau}-\log\tilde{M}$ relation: $\log\bar{\tau}=\gamma(\log\tilde{M}-14)+\log\beta$, in which $\gamma=0.55\pm 0.1$. We also vary the aperture photometric filter radius and measure the $\bar{\tau}$ profiles of group samples, whose constraints on the baryon distribution within and around dark matter halos will be discussed in a companion paper.
@@ 1 Introduction
The kinetic Sunyaev Zel’dovich (kSZ) effect is a secondary cosmic microwave background (CMB) effect induced by the scattering of the CMB photon off free electrons with a bulk motion (Sunyaev & Zeldovich, 1970, 1972, 1980; Phillips, 1995; Birkinshaw, 1999). Its detection enables detailed studies of the baryon distribution in the universe and the cosmological model. On one hand, the kSZ signal is linearly proportional to the free electron number density and is independent of the gas temperature. It can be used to constrain the Epoch-of-Reionization (EoR) history (Alvarez, 2016; Chen et al., 2023b) or the gas distribution within halos (Schaan et al., 2016; Sugiyama et al., 2018a; Schneider et al., 2022) and filaments (Zheng et al., 2023), and is well suited to search for missing baryons which are believed to mainly reside in low density and low temperature environments. On the other hand, the kSZ effect captures the peculiar velocity field of our universe and records its structure growth information. It can help constrain different aspects of the cosmological model such as Copernican principle (Zhang & Stebbins, 2011), primordial non-Gaussianity (Münchmeyer et al., 2019; Kumar et al., 2022), dark energy (Pen & Zhang, 2014; Okumura & Taruya, 2022), modified gravity (Roncarelli et al., 2018; Zheng, 2020; Mitchell et al., 2021), massive neutrinos (Roncarelli et al., 2017) and so on.
@@
Albeit being rich in cosmological information, the detection of the kSZ effect is difficult in its nature. The amplitude of the kSZ signal is two orders of magnitudes lower than the CMB primary fluctuation and several times lower than the thermal Sunyaev Zel’dovich (tSZ) signal (Zhang et al., 2004). Besides, its spectrum is the same as that of the primary CMB, thus multi-frequency measurements do not help on its detection, in contrast to the tSZ effect. As a result, the current detection signal-to-noise (S/N) of the separated kSZ component from the CMB power spectrum is still low (Dunkley et al., 2013; George et al., 2015; Bleem et al., 2022). Fortunately, galaxy resides in high electron density environment and its celestial position denotes where the kSZ signal is relatively high. By cross-correlating the CMB data with the galaxy redshift survey data, we up-weight the high kSZ signal regions and enhance the measurement S/N (Hand et al., 2012). Adopting this so-called kSZ tomography technique, we have detected the late-time (post Epoch-of-Reionization) kSZ effect in a significance around $4\sim 7\sigma$, with combinations of different CMB data and galaxy catalogs, e.g., SPT+DES (Soergel et al., 2016), ACT+SDSS (Schaan et al., 2021; Calafut et al., 2021), Planck/WMAP+unWISE (Kusiak et al., 2021), Planck+DESI(photo-z) (Chen et al., 2022), et al. With the next generation of surveys, S/Ns of such measurements can reach $\mathcal{O}(100)$ and beyond (Sugiyama et al., 2018a; Smith et al., 2018).
@@
The kSZ tomography technique can by implemented by various estimators. Most of them are mathematically equivalent to the three-point correlation function (or bispectrum) of the type $\langle ggT\rangle$ or $\langle gTT\rangle$ (Smith et al., 2018). Here ‘$g$’ is the power of the galaxy field and ‘$T$’ is the power of the CMB field. The majority of such estimators are $\langle ggT\rangle$-like, such as (1) the pairwise estimator, in which we sum the pair difference of the high-pass filtered CMB at galaxy positions within a certain pair separation bin (Hand et al., 2012; Sugiyama et al., 2018a; Gong et al., 2023); (2) the kSZ template method, in which the kSZ template is cross-correlated with the CMB map to extract the kSZ signal, and the core of the kSZ template is the projected galaxy momentum field which is reconstructed from the observed galaxy density field (Ho et al., 2009; Shao et al., 2011); (3) the stacked velocity matched filter method, in which we stack all the high-pass filtered CMB at galaxy locations weighted by the galaxies’ velocities to detect the kSZ signal, and the galaxy velocity filed is reconstructed from the galaxy density field (Li et al., 2014; Schaan et al., 2016, 2021); (4) the growth reconstructed method, in which we directly compare the velocity field constructed from the galaxy density field with the high-pass filtered CMB data at galaxy locations with the Maximum Likelihood Estimation method and extract the structure growth information (Alonso et al., 2016); (5) the velocity reconstruction method, in which we directly reconstruct the large scale radial velocity field from the galaxy density weighted CMB data with a quadratic estimator (Deutsch et al., 2018) or via the machine learning technique (Wang et al., 2021). The $\langle gTT\rangle$-like estimator includes the one by cross-correlating the large scale structure with the squared high-pass filtered CMB data (DeDeo et al., 2005; Hill et al., 2016; Ferraro et al., 2016; Patki et al., 2023). Besides these three-point ones, a higher order estimator for the kSZ tomography with the $\langle gggT\rangle$-type has also been proposed (Kuruvilla, 2022).
@@
A dual analysis both in configuration and Fourier space has been a common practice in the galaxy clustering study. Although the correlation function and the power spectrum are Fourier counterparts and contain the same cosmological information in an ideal case, the cosmic variance on large scales, the shot noise on small scales, and different observational and theoretical systematics restrict the available data within a certain scale range and render different pros and cons to these two statistics (Verde, 2007). For example, the correlation function is straightforward to detect observationally while its covariance matrix estimation and redshift space modelling is non-trivial. On the opposite, the power spectrum has a more diagonal covariance matrix and its anisotropic clustering property is easier to model in redshift space, yet its measurement is contaminated by more systematics such as the anisotropic survey window function and the complicated selection function. In this sense, analysis in both spaces are complementary to each other and are both valuable in constraining cosmological models. Following this spirit, we implement a Fourier space kSZ analysis in this work, while a complementary configuration space kSZ detection from the similar data set has been published in Chen et al. (2022).
@@
The first and only Fourier space kSZ detection until now is from Sugiyama et al. (2018a). We follow the main routine developed in that paper and make several modifications/improvements, including: (1) the median group masses ($\tilde{M}$) of our samples range from $10^{13}-10^{14}{\rm M}_{\odot}/h$, in which the heaviest $\tilde{M}\sim 10^{14}{\rm M}_{\odot}/h$ is larger than those of most other works, and by this we explore a new range of halo mass in the kSZ detection (see also Soergel et al. (2016)); (2) we implement the aperture photometry (AP) filter in the spherical harmonic space (Chen et al., 2022), which can avoid the ambiguity of the pixelized temperature average on the circular filter in configuration space; (3) when applying the kSZ power spectrum estimator including the moving line-of-sight (LOS) effect, we decompose the Legendre polynomial into into a product of spherical harmonics (Hand et al., 2017; Sugiyama et al., 2018b), instead of a Cartesian decomposition (Bianchi et al., 2015; Scoccimarro, 2015), and this will reduce the number of needed FFTs if we include the multipole calculation higher than the dipole in the future; (4) instead of the linear model, we apply a non-linear kSZ model developed in Zheng (2020); Xiao & Zheng (2023) to fit the measured kSZ power spectrum, and since the galaxy group data is from the photometric observation, we further include the photo-z effect in the model; (5) we measure and analyze the redshift space density power spectrum from the same group catalog to obtain constraints on the galaxy bias parameters $b_{1}$, $b_{2}$ and photo-z error dispersion $\sigma_{\rm pho-z}$, by which we break the $b_{1}-\bar{\tau}$ degeneracy in a self-consistent way and avoid possible systematic errors coming from manually selected $b_{1}-M$ relation and $\sigma_{\rm pho-z}$.
@@
The paper is organized as follows. In section 2, we introduce the CMB map and the galaxy group/cluster catalogs we will analyze in this work. In section 3, a kSZ power spectrum model taking into account the photo-z error is developed. In section 4, we outline the methodology we imply to measure the kSZ power spectrum, and in section 5, we display the main results of this work. We conclude in section 6 and more technique details and complementary results are shown in appendices. We will adopt the Planck2018 cosmology (Planck Collaboration et al., 2020) throughout the work, and calculations of cosmological quantities are made via the Nbodykit python toolkit [^1] https://github.com/bccp/nbodykit (Hand et al., 2018).
@@ 2 Data
The detection of the pairwise kSZ effect requires the synergy between a CMB map and a galaxy (group/cluster) catalog. The angular position of a galaxy or galaxy group indicates a sky location likely with a high kSZ signal. A compensated AP filter is applied to the CMB data on this location, filtering out the contamination from the primordial CMB fluctuation and the detector noise, and extract the desired kSZ signal. All these individual kSZ signals are then fed into the pairwise kSZ estimator (eqs. (4) and (5)) and the output is the statistical pairwise kSZ effect. This pairwise property of the estimator tackles the directional dependence of the kSZ signal since the result simply vanishes in a trivial average of signals from all locations due to the isotropic peculiar velocity field. In this section we first introduce the data sets that we are gonna analyze in this work.
@@ 2.1 Planck data
Figure 1: The sky distribution of groups (left) and random points (right) of the baseline sample shown in the Galactic coordinate with $N_{\rm side}=256$ in the HEALPix grid frame. The colour bar represents the number of groups or random points in each pixel. There are $\sim 5.65\times 10^{6}$ groups in the baseline sample, and the random catalog size is 20 times of the corresponding group sample.
The kSZ signal is measured on a full-sky intensity map, HFI 217 GHz [^2] https://irsa.ipac.caltech.edu/data/Planck/release_3/all-sky-maps/ of the public Planck Release 3 data [^3] Based on observations obtained with Planck(http://www.esa.int/Planck), an ESA science mission with instruments and contributions directly funded by ESA Member States, NASA, and Canada.. This map has an effective FWHM $\approx$ 4.87 arcmin and is provided in the HEALPix grid frame (Górski et al., 2005) with $N_{\rm side}=2048$. We update the map to $N_{\rm side}=8096$ with the zeroth-order interpolation in order to improve the map resolution when applying the AP filter (section 4.1).
The contamination of the tSZ effect is largely alleviated due to its vanishment in the frequency of 217 GHz. Although its residual is nonzero due to the finite bandwidth, it is not expected to have the directional dependence as the pairwise kSZ signal does, thus will be further eliminated in our estimator and does not bias the signal. The same cleaning procedure happens for different foregrounds on the CMB map when we apply the pairwise estimator. Compared to the CMB foreground cleaned maps such as SMICA, SEVEM, NILC and COMMANDER, the 217 GHz map has higher angular resolution, which is highly desired in the kSZ detection. By making comparison between different maps in the Appendix C of Chen et al. (2022), the 217 GHz map is shown to be optimal in our kSZ measurement.
@@ 2.2 DESI photometric Data Release 9 galaxy groups
We use the galaxy group catalog [^4] https://gax.sjtu.edu.cn/data/DESI.html provided by Yang et al. (2021) to identify locations where kSZ signals are large. This catalog is constructed from the Data Released 9 (DR9) of the the DESI Legacy Imaging Survey (Dey et al., 2019), by extending the halo-based group finder developed in Yang et al. (2005). After removing the sky area within $|b|<25.0^{\circ}$ and around nearby sources or masked pixels, the sky coverage of the catalog remains 9622 $\rm deg^{2}$ in the North galactic cap (NGC) and 8601 $\rm deg^{2}$ in the South galactic cap (SGC). In total, the group catalog contains 51.44 million galaxy groups in NGC and 45.41 million galaxy groups in SGC with 3D coordinate, richness, halo mass and the total group luminosity.
@@
The halo gas mass is expected to increase with its total mass (e.g., Chen et al. (2022)), therefore we construct group catalogs by varying a lower group mass threshold $M_{\rm th}$. From this, we construct group samples with median masses ($\tilde{M}$) ranging from $10^{13}-10^{14}{\rm M}_{\odot}/h$. In appendix D we show how the detection S/N varies with $M_{\rm th}$, and we recognize the group sample with the highest detection S/N as the baseline sample in this work.
@@
The baseline sample we construct contains $\sim 5.65\times 10^{6}$ heaviest groups in the group catalog, spanning between $0.920<z<$ 0.92555We truncate the redshift distribution at $z=0.92$ to avoid a spurious redshift distribution peak at $z=0.93$ due to some bugs in the DESI spectroscopic redshifts of the current version.. Their pixelized sky distribution is shown on the left panel of figure 1, and their mass, redshift, angular radius and richness distributions are shown in figure 2. We mark the median values of $\{M,z,\theta_{\rm halo},\rm richness\}$ for the NGC and SGC sub-baseline samples in figure 2. The $\{M,z\}$ medians will be used in our following calculation of the theoretical kSZ power spectrum. In Section 5 we will present the measured kSZ power spectum and associated results from the baseline sample.
@@
On the lower left panel of figure 2, half groups in the baseline sample have an angular radius lower than $1.5^{\prime}$, much smaller than the Planck FWHM, which results in a heavy smoothing effect on the kSZ signal and limits its S/N. In this sense, the Planck data is not ideal for the SZ effect detection, although its all-sky coverage compensates this disadvantage. On the contrary, high resolution CMB experiments such as ACT and SPT are limited by their relatively small sky coverage overlaps with current galaxy surveys (Calafut et al., 2021; Bleem et al., 2022). On the lower right panel of figure 2, half groups contain $\leq 3$ group members, and most of them reside at high redshifts. The typical photo-z error of the DESI photometric galaxies can be parameterized by $\sigma_{\rm pho-z}=0.01+0.015z$ (Zhou et al., 2021; Wang et al., 2020; Yang et al., 2021). For groups with richness $\geq 3$, Yang et al. (2021) identified that their $\sigma_{\rm pho-z}$ ranges from 0.005 at $z=0$ to 0.015 at $z=1$. These priors indicate that our group samples will have typical photo-z errors of $\sigma_{\rm pho-z}=0.005\sim 0.025$. Such large redshift uncertainty will heavily dilute group clustering along the line-of-sight. Moreover, regarding that our selected samples span large redshift ranges, the complicated photo-z probability distribution function (PDF), which is redshift dependent and possibly asymmetric, makes the modelling of the galaxy group kSZ and density power spectra in redshift space a challenging task. In appendix A we test the sensitivity of the optical depth measurement on different photo-z damping models and make sure our results are not significantly biased by the photo-z uncertainty.

### 01 Mar 2024 v2

# Detection of pairwise kinetic Sunyaev–Zel’dovich effect with DESI galaxy groups and Planck in Fourier space

@@ Abstract
We report a $\sim 5.2\sigma$ detection of the kinetic Sunyaev–Zel’dovich (kSZ) effect in Fourier space, by combining the DESI galaxy groups and the Planck data. We use the density-weighted pairwise kSZ power spectrum as the summary statistic, and the detailed procedure of its measurement is presented in this paper. Meanwhile, we analyze the redshift space group density power spectrum to constrain its bias parameters and photo-z uncertainties. These best-fitted parameters are substituted to a nonlinear kSZ model, and we fit the measured kSZ power spectrum with this model to constrain the group optical depth $\bar{\tau}$. Selected by a varying lower mass threshold $M_{\rm th}$, the galaxy group catalogs with different median masses ($\tilde{M}$) are constructed from the DR9 data of the DESI Legacy Imaging Surveys. $\tilde{M}$ spans a wide range of $\sim 10^{13}-10^{14}M_{\odot}/h$ and the heaviest $\tilde{M}\sim 10^{14}M_{\odot}/h$ is larger than those of most other kSZ detections. When the aperture photometric filter radius $\theta_{\rm AP}$ is set to be $4.2$ arcmin, the $\tilde{M}=1.75\times 10^{13}M_{\odot}/h$ group sample at the median redshift $\tilde{z}=0.64$ has the highest kSZ detection ${\rm S/N}=5.2$. By fitting $\bar{\tau}$s from various samples against their $\tilde{M}$s, we obtain a linear $\log\bar{\tau}-\log\tilde{M}$ relation: $\log\bar{\tau}=\gamma(\log\tilde{M}-14)+\log\beta$, in which $\gamma=0.55\pm 0.1$. We also vary the aperture photometric filter radius and measure the $\bar{\tau}$ profiles of group samples, whose constraints on the baryon distribution within and around dark matter halos will be discussed in a companion paper.
@@ 1 Introduction
The kinetic Sunyaev-Zel’dovich (kSZ) effect is a secondary cosmic microwave background (CMB) effect induced by the scattering of the CMB photon off free electrons with a bulk motion (Sunyaev & Zeldovich, 1970, 1972, 1980; Phillips, 1995; Birkinshaw, 1999). Its detection enables detailed studies of the baryon distribution in the universe and the cosmological model. On one hand, the kSZ signal is linearly proportional to the free electron number density and is independent of the gas temperature. It can be used to constrain the Epoch-of-Reionization (EoR) history (Alvarez, 2016; Chen et al., 2023b) or the gas distribution within halos (Schaan et al., 2016; Sugiyama et al., 2018a; Chaves-Montero et al., 2021; Schneider et al., 2022) and filaments (Zheng et al., 2023), and is well suited to search for missing baryons which are believed to mainly reside in low density and low temperature environments. On the other hand, the kSZ effect captures the peculiar velocity field of our universe and records its structure growth information. It can help constrain different aspects of the cosmological model such as Copernican principle (Zhang & Stebbins, 2011), primordial non-Gaussianity (Münchmeyer et al., 2019; Kumar et al., 2022), dark energy (Pen & Zhang, 2014; Okumura & Taruya, 2022), modified gravity (Bianchini & Silvestri, 2016; Roncarelli et al., 2018; Zheng, 2020; Mitchell et al., 2021), massive neutrinos (Roncarelli et al., 2017) and so on.
@@
Albeit, being rich in cosmological information, the detection of the kSZ effect is difficult in its nature. The amplitude of the kSZ signal is 2 orders of magnitudes lower than the CMB primary fluctuation and several times lower than the thermal Sunyaev-Zel’dovich (tSZ) signal (Zhang et al., 2004). Besides, its spectrum is the same as that of the primary CMB; thus, multifrequency measurements do not help on its detection, in contrast to the tSZ effect. As a result, the current detection signal-to-noise ratio (S/N) of the separated kSZ component from the CMB power spectrum is still low (Dunkley et al., 2013; George et al., 2015; Bleem et al., 2022). Fortunately, the galaxy resides in high electron density environment and its celestial position denotes where the kSZ signal is relatively high. By cross-correlating the CMB data with the galaxy redshift survey data, we up-weight the high kSZ signal regions and enhance the measurement S/N (Hand et al., 2012). Adopting this so-called kSZ tomography technique, we have detected the late-time (post Epoch-of-Reionization) kSZ effect in a significance around $4\sigma\sim 7\sigma$, with combinations of different CMB data and galaxy catalogs, e.g., South Pole Telescope (SPT)+Dark Energy Survey (DES) (Soergel et al., 2016; Schiappucci et al., 2023), Atacama Cosmology Telescope (ACT)+SDSS (Schaan et al., 2021; Calafut et al., 2021), Planck/Wilkinson Microwave Anisotropy Probe+unWISE (Kusiak et al., 2021), Planck+DESI(photo-z) (Chen et al., 2022), et al. With the next generation of surveys, S/Ns of such measurements can reach $\mathcal{O}(100)$ and beyond (Sugiyama et al., 2018a; Smith et al., 2018).
@@
The kSZ tomography technique can by implemented by various estimators. Most of them are mathematically equivalent to the three-point correlation function (or bispectrum) of the type $\langle ggT\rangle$ or $\langle gTT\rangle$ (Smith et al., 2018). Here, “$g$” is the power of the galaxy field, and “$T$” is the power of the CMB field. The majority of such estimators are $\langle ggT\rangle$-like, such as (1) the pairwise estimator, in which we sum the pair difference of the high-pass filtered CMB at galaxy positions within a certain pair separation bin (Hand et al., 2012; Sugiyama et al., 2018a; Gong et al., 2023); (2) the kSZ template method, in which the kSZ template is cross-correlated with the CMB map to extract the kSZ signal, and the core of the kSZ template is the projected galaxy momentum field which is reconstructed from the observed galaxy density field (Ho et al., 2009; Shao et al., 2011); (3) the stacked velocity matched filter method, in which we stack all the high-pass filtered CMB at galaxy locations weighted by the galaxies’ velocities to detect the kSZ signal, and the galaxy velocity filed is reconstructed from the galaxy density field (Li et al., 2014; Schaan et al., 2016, 2021); (4) the growth reconstructed method, in which we directly compare the velocity field constructed from the galaxy density field with the high-pass filtered CMB data at galaxy locations with the maximum likelihood estimation method and extract the structure growth information (Alonso et al., 2016); (5) the velocity reconstruction method, in which we directly reconstruct the large-scale radial velocity field from the galaxy density-weighted CMB data with a quadratic estimator (Deutsch et al., 2018) or via the machine-learning technique (Wang et al., 2021). The $\langle gTT\rangle$-like estimator includes the one by cross-correlating the large-scale structure with the squared high-pass filtered CMB data (DeDeo et al., 2005; Hill et al., 2016; Ferraro et al., 2016; Patki et al., 2023). Besides these three-point ones, higher order estimators for the kSZ tomography with the $\langle gggT\rangle$-type has also been proposed (Hernández-Monteagudo et al., 2020; Chaves-Montero et al., 2021; Kuruvilla, 2022).
@@
A dual analysis both in configuration and Fourier space has been a common practice in the galaxy clustering study. Although the correlation function and the power spectrum are Fourier counterparts and contain the same cosmological information in an ideal case, the cosmic variance on large-scales, the shot noise on small scales, and different observational and theoretical systematics restrict the available data within a certain scale range and render different pros and cons to these two statistics (Verde, 2007). For example, the correlation function is straightforward to detect observationally while its covariance matrix estimation and redshift space modeling is nontrivial. On the other hand, the power spectrum has a more diagonal covariance matrix and its anisotropic clustering property is easier to model in redshift space, yet its measurement is contaminated by more systematics such as the anisotropic survey window function and the complicated selection function. In this sense, the analysis in both spaces are complementary to each other and are both valuable in constraining cosmological models. Following this spirit, we implement a Fourier space kSZ analysis in this work, while a complementary configuration space kSZ detection from the similar data set has been published in Chen et al. (2022).
@@
The first and only Fourier space kSZ detection until now is from Sugiyama et al. (2018a). We follow the main routine developed in that paper and make several modifications/improvements, including the following: (1) the median group masses ($\tilde{M}$) of our samples range from $10^{13}-10^{14}M_{\odot}/h$, in which the heaviest $\tilde{M}\sim 10^{14}M_{\odot}/h$ is larger than those of most other works, and by this we explore a new range of halo mass in the kSZ detection (see also Soergel et al. (2016)); (2) we implement the aperture photometry (AP) filter in the spherical harmonic space (Chen et al., 2022), which can avoid the ambiguity of the pixelized temperature average on the circular filter in configuration space; (3) when applying the kSZ power spectrum estimator including the moving line-of-sight (LOS) effect, we decompose the Legendre polynomial into into a product of spherical harmonics (Hand et al., 2017; Sugiyama et al., 2018b), instead of a Cartesian decomposition (Bianchi et al., 2015; Scoccimarro, 2015), and this will reduce the number of needed fast Fouirer transforms (FFTs) if we include the multipole calculation higher than the dipole in the future; (4) instead of the linear model, we apply a nonlinear kSZ model developed in Zheng (2020); Xiao & Zheng (2023) to fit the measured kSZ power spectrum, and since the galaxy group data are from the photometric observation, we further include the photo-z effect in the model; (5) we measure and analyze the redshift space density power spectrum from the same group catalog to obtain constraints on the galaxy bias parameters $b_{1}$, $b_{2}$ and photo-z error dispersion $\sigma_{\rm pho-z}$, by which we break the $b_{1}-\bar{\tau}$ degeneracy in a self-consistent way and avoid possible systematic errors coming from manually selected $b_{1}-M$ relation and $\sigma_{\rm pho-z}$.
@@
The paper is organized as follows. In Section 2, we introduce the CMB map and the galaxy group/cluster catalogs we will analyze in this work. In section 3, a kSZ power spectrum model taking into account the photo-z error is developed. In Section 4, we outline the methodology we imply to measure the kSZ power spectrum, and in Section 5, we display the main results of this work. We conclude in Section 6 and more technique details and complementary results are shown in appendices. We will adopt the Planck2018 cosmology (Planck Collaboration et al., 2020) throughout the work, and calculations of cosmological quantities are made via the Nbodykit Python toolkit [^1] https://github.com/bccp/nbodykit from Hand et al. (2018).
@@ 2 Data
The detection of the pairwise kSZ effect requires the synergy between a CMB map and a galaxy (group/cluster) catalog. The angular position of a galaxy or galaxy group indicates a sky location likely with a high kSZ signal. A compensated AP filter is applied to the CMB data on this location, filtering out the contamination from the primordial CMB fluctuation and the detector noise, and extracting the desired kSZ signal. All these individual kSZ signals are then fed into the pairwise kSZ estimator (Equations (4) and (5)) and the output is the statistical pairwise kSZ effect. This pairwise property of the estimator tackles the directional dependence of the kSZ signal since the result simply vanishes in a trivial average of signals from all locations due to the isotropic peculiar velocity field. In this section we first introduce the data sets that we are going to analyze in this work.
@@ 2.1 Planck data
Figure 1: The sky distribution of groups (left) and random points (right) of the baseline sample shown in the Galactic coordinate with $N_{\rm side}=256$ in the HEALPix grid frame. The color bar represents the number of groups or random points in each pixel. There are $\sim 5.65\times 10^{6}$ groups in the baseline sample, and the random catalog size is 20 times of the corresponding group sample.
The kSZ signal is measured on a full-sky intensity map, High Frequency Instrument (HFI) 217 GHz [^2] https://irsa.ipac.caltech.edu/data/Planck/release_3/all-sky-maps/ of the public Planck Release 3 data [^3] Based on observations obtained with Planck(http://www.esa.int/Planck), an ESA science mission with instruments and contributions directly funded by ESA Member States, NASA, and Canada.. This map has an effective FWHM $\approx$ 4.87 arcmin and is provided in the HEALPix grid frame (Górski et al., 2005) with $N_{\rm side}=2048$. We update the map to $N_{\rm side}=8096$ with the zeroth-order interpolation in order to improve the map resolution when applying the AP filter (Section 4.1).
The contamination of the tSZ effect is largely alleviated due to its vanishment in the frequency of 217 GHz. Although its residual is nonzero due to the finite bandwidth, it is not expected to have the directional dependence as the pairwise kSZ signal does, thus will be further eliminated in our estimator and does not bias the signal. The same cleaning procedure happens for different foregrounds on the CMB map when we apply the pairwise estimator. Compared to the CMB foreground cleaned maps such as SMICA, SEVEM, NILC and COMMANDER, the 217 GHz map has higher angular resolution, which is highly desired in the kSZ detection. By making comparison between different maps in Appendix C of Chen et al. (2022), the 217 GHz map is shown to be optimal in our kSZ measurement.
@@ 2.2 DESI photometric Data Release 9 galaxy groups
We use the galaxy group catalog [^4] https://gax.sjtu.edu.cn/data/DESI.html provided by Yang et al. (2021) to identify locations where kSZ signals are large. This catalog is constructed from the Data Release 9 (DR9) of the the DESI Legacy Imaging Survey (Dey et al., 2019), by extending the halo-based group finder developed in Yang et al. (2005). After removing the sky area within $|b|<25.^{\circ}0$ and around nearby sources or masked pixels, the sky coverage of the catalog remains 9622 $\rm deg^{2}$ in the North galactic cap (NGC) and 8601 $\rm deg^{2}$ in the South galactic cap (SGC). In total, the group catalog contains 51.44 million galaxy groups in NGC and 45.41 million galaxy groups in SGC with 3D coordinate, richness, halo mass and the total group luminosity.
@@
The halo gas mass is expected to increase with its total mass (e.g., Chen et al. (2022)); therefore, we construct group catalogs by varying a lower group mass threshold $M_{\rm th}$. From this, we construct group samples with median masses ($\tilde{M}$) ranging from $10^{13}-10^{14}M_{\odot}/h$. In Appendix D, we show how the detection S/N varies with $M_{\rm th}$, and we recognize the group sample with the highest detection S/N as the baseline sample in this work.
@@
The baseline sample we construct contains $\sim 5.65\times 10^{6}$ heaviest groups in the group catalog, spanning between $0.920<z<$ 0.92555We truncate the redshift distribution at $z=0.92$ to avoid a spurious redshift distribution peak at $z=0.93$ due to some bugs in the DESI spectroscopic redshifts of the current version.. Their pixelized sky distribution is shown on the left panel of figure 1, and their mass, redshift, angular radius and richness distributions are shown in figure 2. We mark the median values of $\{M,z,\theta_{\rm halo},\rm richness\}$ for the NGC and SGC subbaseline samples in figure 2. The $\{M,z\}$ medians will be used in our following calculation of the theoretical kSZ power spectrum. In Section 5, we will present the measured kSZ power spectum and associated results from the baseline sample.
@@
On the lower left panel of figure 2, half of the groups in the baseline sample have an angular radius lower than $1.5^{\prime}$, much smaller than the Planck FWHM, which results in a heavy smoothing effect on the kSZ signal and limits its S/N. In this sense, the Planck data are not ideal for the SZ effect detection, although its all-sky coverage compensates this disadvantage. On the contrary, high resolution CMB experiments such as ACT and SPT are limited by their relatively small sky coverage overlaps with current galaxy surveys (Calafut et al., 2021; Bleem et al., 2022). On the lower right panel of figure 2, half of the groups contain $\leq 3$ group members, and most of them reside at high redshifts. The typical photo-z error of the DESI photometric galaxies can be parameterized by $\sigma_{\rm pho-z}=0.01+0.015z$ (Zhou et al., 2021; Wang et al., 2020; Yang et al., 2021). For groups with richness $\geq 3$, Yang et al. (2021) identified that their $\sigma_{\rm pho-z}$ ranges from 0.005 at $z=0$ to 0.015 at $z=1$. These priors indicate that our group samples will have typical photo-z errors of $\sigma_{\rm pho-z}=0.005\sim 0.025$. Such large redshift uncertainty will heavily dilute group clustering along the LOS. Moreover, regarding that our selected samples span large redshift ranges, the complicated photo-z probability distribution function (PDF), which is redshift dependent and possibly asymmetric, makes the modeling of the galaxy group kSZ and density power spectra in redshift space a challenging task. In Appendix A we test the sensitivity of the optical depth measurement on different photo-z damping models and make sure our results are not significantly biased by the photo-z uncertainty.